### v1

# RespoDiff: Dual-Module Bottleneck Transformation for Responsible & Faithful T2I Generation

Abstract
The rapid advancement of diffusion models has enabled high-fidelity and semantically rich text-to-image generation; however, ensuring fairness and safety remains an open challenge. Existing methods typically improve fairness and safety at the expense of semantic fidelity and image quality. In this work, we propose RespoDiff, a novel framework for responsible text-to-image generation that incorporates a dual-module transformation on the intermediate bottleneck representations of diffusion models. Our approach introduces two distinct learnable modules: one focused on capturing and enforcing responsible concepts, such as fairness and safety, and the other dedicated to maintaining semantic alignment with neutral prompts. To facilitate the dual learning process, we introduce a novel score-matching objective that enables effective coordination between the modules. Our method outperforms state-of-the-art methods in responsible generation by ensuring semantic alignment while optimizing both objectives without compromising image fidelity. Our approach improves responsible and semantically coherent generation by ~20% across diverse, unseen prompts. Moreover, it integrates seamlessly into large-scale models like SDXL, enhancing fairness and safety. Code will be released upon acceptance.
1 Introduction
The rise of text-to-image diffusion models (T2I), such as Stable Diffusion, has transformed content creation and visual communication by generating high-quality visuals from text prompts (Podell et al., 2024; Rombach et al., 2022). However, these models risk reinforcing stereotypes or producing harmful content, which can lead to societal consequences (Luccioni et al., 2023; Perera and Patel, 2023; Rando et al., 2022; Schramowski et al., 2023). Ensuring a responsible workflow is critical to mitigating these risks.
Previous methods for responsible text-to-image (T2I) generation include prompt modification (Chuang et al., 2023; Ni et al., 2023), model fine-tuning (Gandikota et al., 2023; Shen et al., 2024), model editing (Gandikota et al., 2024), classifier guidance (Schramowski et al., 2023), and latent vector injection (Li et al., 2024). Despite advancements in responsible T2I generation, many existing approaches still compromise semantic fidelity and image quality, reducing their effectiveness in producing both responsible and faithful generation. To address these challenges, we propose RespoDiff, a novel framework for responsible T2I generation that introduces a dual-module transformation on the bottleneck representations of diffusion models. Specifically, given a responsible category, such as demographic attributes (e.g., gender) or safety factors – and its associated target concepts (e.g.,“man”, “woman” etc.), our approach learns independent transformations that steer the diffusion model toward target-aligned outputs while maintaining coherence with a neutral prompt (e.g., “a person”). These learned transformations can then be applied during inference to promote fairness and safety in T2I generation without compromising the underlying structure of the diffusion process. RespoDiff incorporates two learnable modules: ① Responsible Concept Alignment Module, which steers latent representations toward fair and safe outputs by learning transformations that align with responsible target concepts; and ② Semantic Alignment Module, which preserves consistency with neutral prompts, ensuring the generated images remain aligned with original prompts.
At the core of RespoDiff is a score-matching objective that coordinates the two modules. For the demographic category “gender”, given a neutral prompt (e.g., “a person”) and a target concept (e.g., “a woman”), we introduce an objective to guide the Responsible Concept Alignment Module, which learns to modify the diffusion trajectory of the neutral latent such that it closely approximates the trajectory associated with the target concept. A key aspect of our approach is leveraging the neutral denoised latent, obtained by passing the neutral prompt through the diffusion model. This serves as a stable reference point, allowing us to extract explicit directional guidance by comparing UNet predictions for the neutral and target concepts. This ensures that the transformation consistently steers generation toward the desired concept.
Additionally, we introduce a score-matching objective to mitigate excessive influence from the learned transformation and prevent oversteering toward the target concept. This objective updates the Semantic Alignment Module, ensuring that the dual-module transformation stays aligned with the original generative trajectory for the neutral prompt. The Semantic Alignment Module safeguards the structure and semantic details of the image with respect to the neutral prompt, thereby preserving visual fidelity of the original diffusion model. By jointly optimizing these objectives, our method ensures that the generated outputs align with the target concept while maintaining all other visual details consistent with the neutral prompt.
We empirically validate the effectiveness of our approach for responsible T2I generation, with a focus on fairness and safety. RespoDiff surpasses existing fair-generation baselines and effectively generalizes to unseen prompts, including profession-specific scenarios, without requiring any profession-specific training or fine-tuning. Our approach further ensures semantic alignment with prompts while maintaining the visual quality of diffusion models. Additionally, it eliminates harmful or unsafe outputs without compromising image fidelity or alignment, demonstrating the practicality and robustness of our framework. Notably, our framework can seamlessly be integrated into large-scale T2I models such as SDXL, enhancing fairness and safety in real-world deployments.
The key contributions of this work are as follows: We introduce RespoDiff, a novel dual-module transformation for diffusion models, integrating a Responsible Concept Alignment module with a Semantic Alignment Module to ensure responsible generation while being faithful to the original diffusion process. We propose a simple score-matching objective that enables effective coordination between the modules, ensuring seamless integration of responsible generation and prompt alignment. Our method achieves approximately 20% improvement in fairness and safety metrics while ensuring high semantic fidelity and image quality, demonstrating robustness across unseen prompts.
2 Related Work
T2I generation has transformed generative AI, enabling highly realistic image creation from text (Ho et al., 2020a; Ramesh et al., 2022; Rombach et al., 2022), but also raises ethical concerns, including the risk of generating harmful or inappropriate content (Cho et al., 2023; Luccioni et al., 2023).
Responsible Generation using Diffusion Models: In recent years, there has been a growing emphasis on methods to reduce biased and inappropriate content generation in diffusion models, such as Stable Diffusion. Several approaches focus on modifying input prompts by removing harmful or problematic terms (Ni et al., 2023; Schramowski et al., 2023), while others employ prompt-tuning techniques (Kim et al., 2023) or learn projection embeddings on prompt representations (Chuang et al., 2023) to filter out undesirable content. Some methods, such as those by Gandikota et al. (2023), Huang et al. (2023) and Zhang et al. (2024) attempt to erase unsafe concept representations from the diffusion models, though these techniques may negatively affect the model’s original performance. Similarly, Shen et al. (2024) addresses biases by fine-tuning specific parts of the model weights, but such approaches require additional training for each prompt or domain. In contrast, our method avoids prompt-specific fine-tuning and generalizes effectively across diverse and unseen prompts, including profession-specific ones. Alternative strategies such as those by Friedrich et al. (2023), Parihar et al. (2024), and Schramowski et al. (2023) use classifier-free guidance to steer the generation process away from undesirable content without requiring extra training. While some methods propose efficient closed-form solutions for embedding matrices to ensure responsible content generation (Chuang et al., 2023; Gandikota et al., 2024), they lack the adaptability and fine-grained control over image generation offered by our approach.
Concept Discovery in bottleneck layer: Our method shares the goal of learning responsible concept representations in the latent space of diffusion models. Kwon et al. (2023) were among the pioneers to identify the bottleneck layer of U-Net (the $h$-space) as a semantic latent space, demonstrating that interventions within this space lead to semantically meaningful changes in the generated images. Their approach uses off-the-shelf CLIP classifiers to learn disentangled representations in the $h$-space. Li et al. (2024) built on this work by identifying interpretable directions in the latent space for target concepts. Their approach generates target concept images, adds noise, and then denoises them using a neutral prompt and a learnable vector optimized via noise reconstruction. However, it lacks explicit reference to the neutral denoised latent, relying on indirect supervision that may lead to less precise control over the transformation. Additionally, it does not explicitly enforce faithfulness, risking unintended deviations. In contrast, RespoDiff directly models diffusion trajectory shifts using the neutral denoised latent for precise concept learning while integrating a Semantic Alignment Module to ensure coherence with the original prompt.
Figure 1: Illustration of RespoDiff : RespoDiff performs reverse diffusion to timestep $t$ using the prompt “a person”, obtaining latent $\bm{z}_{t,\text{neu}}$ as the neutral denoised latent for all forward processes. Forward diffusion with the RAM and “a person” predicts a neutral score, with mean-squared error between neutral and target scores updating the RAM. To maintain faithfulness to the original diffusion process, forward diffusion with both RAM and SAM generates a neutral score, with mean-squared error between neutral and original scores updating the SAM.
3 Preliminaries
In this section, we provide the necessary background regarding diffusion models and the scoring functions which forms the foundation of our model design.
Diffusion Models: Diffusion models (Ho et al., 2020b; Sohl-Dickstein et al., 2015) are likelihood-based generative models inspired by nonequilibrium thermodynamics (Song and Ermon, 2019). The model learns a denoising process that transforms random noise into samples from original data distribution, $p_{\text{data}}$. The process involves gradually corrupting training data with Gaussian noise in a forward process, where an initial sample $\bm{x}_{0}\sim p_{\text{data}}$ is progressively noised into $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{T}$ through a Markovian process as follows :
$q(\bm{x}_{1:T}|\bm{x}_{0})=\prod\nolimits_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})\textrm{, }$
$q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t}|\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\mathbf{I})\textrm{,}$
where $T$ is the total number of steps (typically 1000), and variance schedule $\beta_{t}$ ensures $q_{T}(\bm{x}_{T})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$.The reverse process learns to reconstruct the original data by reversing this diffusion.
$p_{\bm{\theta}}(\bm{x}_{0:T})=p(\bm{x}_{T})\prod\nolimits_{t=1}^{T}p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})\textrm{,}$
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1}|\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t),\sigma_{t}\mathbf{I})\textrm{,}$
where $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ is parameterized using a noise prediction network $\epsilon_{\theta}(\bm{x}_{t},t)$. After training, generation in diffusion models involves sampling from $p_{\bm{\theta}}(\bm{x}_{0})$, starting with a noise sample $\bm{x}_{T}\sim p(\bm{x}_{T})$ and recovering $\bm{x}_{0}\sim p_{\text{data}}$ using an SDE/ODE solver (e.g., DDIM (Song et al., 2021)). These models learn the transition probabilities $p(\bm{x}_{t-1}|\bm{x}_{t})$, defined as follows:
$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\bm{x}_{t},t))+\sigma_{t}\bm{w}_{t}\textrm{, }\bm{w}_{t}\sim\mathcal{N}(\bm{0},\mathbf{I})\textrm{,}$ (1)
where $\alpha_{t},\bar{\alpha}_{t}$ and $\beta_{t}$ are predetermined noise variances, and $\bm{w}_{t}$ is a time-dependent weighting function.
Diffusion Scoring Function: The noise prediction network $\epsilon_{\theta}(\bm{x}_{t},t)$ iteratively estimates the noise $\epsilon$ to generate $\bm{x}_{0}$ from $\bm{x}_{T}$ and approximates the score function (Ho et al., 2020b; Song et al., 2020), given by $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\approx-\epsilon_{\theta}(\bm{x}_{t},t)/\sigma_{t}$, where $\sigma_{t}$ is the noise level at time step $t$ and $p_{t}$ is the marginal distribution of the samples noised to time $t$. Following the score function direction guides samples back to the data distribution.
T2I Diffusion Models: T2I models generate images conditioned on text, using a UNet-based noise prediction network $\epsilon_{\theta}(\bm{x}_{t},y,t)$, where $y$ is the text prompt. Rombach et al. (2022) use Latent Diffusion Models (LDMs), where diffusion operates in latent space $\bm{z_{t}}$ instead of image space $\bm{x_{t}}$. Image generation starts by sampling latent noise $\bm{z_{T}}$, applying reverse diffusion to obtain $\bm{z_{0}}$, and decoding it with VAE to obtain $\bm{x_{0}}$. Classifier-free guidance (Ho and Salimans, 2022) is used to enhance conditional generation by adjusting the score function as follows.
$\hat{\epsilon}_{\theta}(\bm{z}_{t};y,t)=\epsilon_{\theta}(\bm{z}_{t};y=\varnothing,t)+s\left(\epsilon_{\theta}(\bm{z}_{t};y,t)-\epsilon_{\theta}(\bm{z}_{t};y=\varnothing,t)\right)$ (2)
where $s$ is the guidance scale and $\epsilon_{\theta}(\bm{z}_{t};y=\varnothing,t)$ denotes the unconditional score. The objective to train T2I diffusion models is given by :
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\bm{z},\epsilon,t}\left[\left\|\hat{\epsilon}_{\theta}(\bm{z}_{t};y,t)-\epsilon\right\|^{2}_{2}\right]$ (3)
4 Method
We propose RespoDiff, a framework for responsible T2I generation via dual transformations on the intermediate representations of diffusion models. An overview is shown in Fig.˜1.
4.1 Problem Formulation
We denote the UNet of a pre-trained T2I diffusion model as $f:\mathcal{Z}\times\mathcal{Y}\to\mathcal{Z}$, where $\mathcal{Z}$ represents the latent space and $\mathcal{Y}$ denotes the set of textual prompts. The UNet consists of an encoder $e:\mathcal{Z}\times\mathcal{Y}\to\mathcal{H}$, which maps a latent $\bm{z}$ and a textual prompt $y$ to an intermediate representation $\bm{h}_{y}\in\mathcal{H}$, and a decoder $g:\mathcal{H}\times\mathcal{Y}\to\mathcal{Z}$, which generates an updated latent $\bm{z^{\prime}}$. Formally, the model is expressed as:
$f(\bm{z},y)=g(e(\bm{z},y),y)=g(\bm{h}_{y},y),$
where $\mathcal{H}$ represents the bottleneck space. Since the encoder and decoder always take the same text input in our setting, we represent $g(\bm{h}_{y},y)=g(\bm{h}_{y})$ for simplicity. The final image is obtained by decoding $\bm{z^{\prime}}$ using a VAE decoder.
To ensure responsible T2I image generation, we focus on a set of responsible concept categories, denoted as $\mathcal{C}=\{\mathcal{C}_{\text{gender}},\mathcal{C}_{\text{race}},\mathcal{C}_{\text{safe}}\}$. Each category comprises sensitive concepts, represented as $\mathcal{S}_{c}$: for example, $\mathcal{S}_{\text{gender}}=\{\text{man},\text{woman}\}$, $\mathcal{S}_{\text{race}}=\{\text{black},\text{asian},\text{white}\}$, and $\mathcal{S}_{\text{safe}}=\{\text{violence},\text{nudity}\}$. These categories and concepts are selected in alignment with prior works to address fairness and safety concerns (Gandikota et al., 2023; Li et al., 2024).
We consider a neutral prompt, denoted as $y_{\text{neu}}$, to facilitate responsible generation. For $\mathcal{C}_{\text{gender}}$ and $\mathcal{C}_{\text{race}}$, we use $y_{\text{neu}}=\text{``a person'' }$, which provides a general description of human subjects. For $\mathcal{C}_{\text{safe}}$, we use $y_{\text{neu}}=\text{``a scene''}$, enabling the model to generalize safety transformations across diverse scenarios. Additionally, target prompts corresponding to each concept $s\in\mathcal{S}_{c}$ are denoted as $y_{\text{tar}}^{s}$.
Let $\bm{h}_{\text{neu}}\in\mathcal{H}$ denote the intermediate bottleneck representations of $f$ corresponding to the neutral prompt $y_{\text{neu}}$. RespoDiff aims to modify $\bm{h}_{\text{neu}}$ such that the transformed model:
$\hat{f}(y_{\text{neu}})=g(\mathcal{T}_{\theta}^{s}(\bm{h}_{\text{neu}}))=g(\mathcal{T}_{\theta}^{\text{resp},s}(\bm{h}_{\text{neu}})+\mathcal{T}_{\theta}^{\text{sem},s}(\bm{h}_{\text{neu}}))$
produces images aligned with the target concept $s\in\mathcal{S}_{c}$, while ensuring semantic alignment and visual quality of original diffusion model $f$. The learnable transformation $\mathcal{T}_{\theta}^{s}$ is decomposed into two components where $\mathcal{T}_{\theta}^{\text{resp},s}$ ensures fairness and safety by aligning generated images with the intended target concept $s$, and $\mathcal{T}_{\theta}^{\text{sem},s}$ preserves the semantic content and visual quality of the images.
By selecting general base descriptions, such as $y_{\text{neu}}=\text{``a person''}$ or $y_{\text{neu}}=\text{``a scene''}$, the learned transformations are designed to generalize effectively across diverse human representations and scene contexts. For fairness, $y_{\text{tar}}^{s}$ corresponds to a concept $s\in\mathcal{S}_{\text{gender}}\cup\mathcal{S}_{\text{race}}$. For safety, following Li et al. (2024), we treat $s\in\mathcal{S}_{\text{safe}}$ as a negative concept, where the associated $y_{\text{tar}}^{s}$ serves as a negative prompt. This encourages the model to steer away from unsafe content by contrasting it with a neutral prompt.
4.2 Responsible Concept Alignment Module
In this section, we introduce the Responsible Concept Alignment Module (RAM) $\mathcal{T}_{\theta}^{\text{resp},s}$, which modifies the latent representation $\bm{h}_{\text{neu}}$ of a neutral prompt to align it with a target concept. Given a neutral prompt $y_{\text{neu}}$ and a target concept prompt $y_{\text{tar}}^{s}$, the goal is to adjust the diffusion trajectory such that the neutral latent evolves to reflect the target concept $s$.
To achieve this, we employ a score-matching objective that directly aligns the model’s latent output with the target concept by manipulating the trajectory of the neutral latent. At a randomly selected timestep $t$, we begin by sampling the neutral denoised latent, $\bm{z}_{t,\text{neu}}$, from the diffusion model $\hat{f}$ through reverse diffusion. This denoised latent becomes the starting point for its transformation towards the target concept. Subsequently, we compute the UNet predictions corresponding to the neutral prompt $y_{\text{neu}}$, denoted as $\epsilon_{\text{neu}}=\epsilon_{f_{\text{resp}}}(\bm{z}_{t,\text{neu}},y_{\text{neu}})$ where $f_{\text{resp}}=g(\mathcal{T}_{\theta}^{\text{resp},s}(\bm{\hat{h}}_{\text{neu}}))$ and $\bm{\hat{h}}_{\text{neu}}=e(\bm{z}_{t,\text{neu}},y_{\text{neu}})$. Using the same denoised latent $\bm{z}_{t,\text{neu}}$, we compute the UNet predictions conditioned on the target prompt $y_{\text{tar}}^{s}$, represented as $\epsilon_{\text{tar}}=\epsilon_{f}(\bm{z}_{t,\text{neu}},y_{\text{tar}}^{s})$, where $f=g(\bm{h}_{\text{tar}})$ and $\bm{h}_{\text{tar}}=e(\bm{z}_{t,\text{neu}},y_{\text{tar}}^{s})$, which serves as the ground truth predictions. The objective is formulated as:
$\mathcal{L}_{\text{resp}}=\mathbb{E}_{\bm{z}_{t,\text{neu}}}\left[\left\|\epsilon_{f_{\text{resp}}}(\bm{z}_{t,\text{neu}},y_{\text{neu}})-\epsilon_{f}(\bm{z}_{t,\text{neu}},y_{\text{tar}}^{s})\right\|_{2}^{2}\right]$ (4)
It is important to note that during this stage, the loss computation considers only $f_{\text{resp}}$, and not $\hat{f}$, as the sole objective is to steer the RAM towards aligning with the target concept. Following Song et al. (2020), the objective in Eq.˜4 can be viewed as minimizing the gradient disparity between the neutral and target diffusion trajectories, effectively guiding the latent space toward the target concept at each timestep. As a result, optimizing this score-matching loss enables the RAM to transform the latent representation such that the generated images are effectively aligned with the desired target concept. Unlike prior methods that rely on implicit reconstructions using pre-existing target images, our loss function explicitly supervises the transformation in the model’s latent space by precisely aligning the diffusion trajectories of neutral and target concepts. By leveraging the neutral denoised latent as a stable reference point, our approach ensures that the modification is structurally guided rather than inferred indirectly from external data.
4.3 Semantic Alignment Module
In this section, we introduce the Semantic Alignment Module (SAM) $\mathcal{T}_{\theta}^{\text{sem},s}$, which ensures that the transformations applied to the latent representations preserve semantic fidelity with respect to the original pre-trained diffusion model $f$. While RAM focuses on aligning with a target concept $s$, the Semantic Alignment Module maintains consistency with the original diffusion trajectory.
To achieve this, we propose a score-matching objective that regularizes the transformation by preserving alignment with the original generative trajectory of a neutral prompt. Specifically, we utilize the denoised latent representation $\bm{z}_{t,\text{neu}}$ obtained from the reverse diffusion process $\hat{f}$ at a randomly selected timestep $t$. Using the default pre-trained diffusion model $f$, we compute the UNet prediction corresponding to the neutral prompt $y_{\text{neu}}$, denoted as $\epsilon_{\text{true}}=\epsilon_{f}(\bm{z}_{t,\text{neu}},y_{\text{neu}})$ where $f=g(\bm{h}_{\text{neu}})$ and $\bm{h}_{\text{neu}}=e(\bm{z}_{t,\text{neu}},y_{\text{neu}})$. This prediction serves as a reference for the original diffusion process. We then introduce a score-matching objective formulated as:
$\mathcal{L}_{\text{sem}}=\mathbb{E}_{\bm{z}_{t,\text{neu}}}\left[\left\|\epsilon_{\hat{f}}(\bm{z}_{t,\text{neu}},y_{\text{neu}})-\epsilon_{f}(\bm{z}_{t,\text{neu}},y_{\text{neu}})\right\|_{2}^{2}\right]$ (5)
This loss penalizes any significant divergence between the steering of the Responsible Concept Alignment Module and the original diffusion model’s behavior. By optimizing $\mathcal{L}_{\text{sem}}$, the Semantic Alignment Module ensures that the generative process adheres closely to the original pre-trained model’s trajectory, thereby preserving semantic fidelity and preventing the introduction of artifacts or unintended deviations.
4.4 Training & inference
The training process alternates between optimizing the transformations $\mathcal{T}_{\theta}^{\text{resp},s}$ and $\mathcal{T}_{\theta}^{\text{sem},s}$ to achieve responsible image generation while maintaining semantic fidelity. We do not backpropagate through the reverse diffusion when obtaining $\bm{z}_{t,\text{neu}}$, keeping the modules frozen primarily to avoid the computational overhead associated with it. In the first step, we update $\mathcal{T}_{\theta}^{\text{resp},s}$ by using the model $f_{\text{resp}}=g(\mathcal{T}_{\theta}^{\text{resp},s}(\bm{\hat{h}}_{\text{neu}}))$ and minimizing the responsible loss $\mathcal{L}_{\text{resp}}$. This update steers the generated images towards responsible concepts while maintaining consistency with the original diffusion process. In the second step, we switch to the full transformation $\hat{f}=g(\mathcal{T}_{\theta}^{\text{resp},s}(\bm{\hat{h}}_{\text{neu}})+\mathcal{T}_{\theta}^{\text{sem},s}(\bm{\hat{h}}_{\text{neu}}))$ and optimize $\mathcal{T}_{\theta}^{\text{sem},s}$ alone using the semantic loss $\mathcal{L}_{\text{sem}}$ weighted by $\lambda$, where $\lambda$ is a hyperparameter. As the iterative optimization progresses, the neutral denoised latent $\bm{z}_{t,\text{neu}}$, obtained via reverse diffusion through $\hat{f}$ with learnable transformations, progressively aligns with both the target concept and the original diffusion process, leading to convergence.
During inference, for responsible generation with respect to a category $c\in\mathcal{C}$, we learn transformations $\mathcal{T}_{\theta}^{s}$ corresponding to each sensitive concept $s\in\mathcal{S}_{c}$. To ensure fairness, we randomly select a concept $s\in\mathcal{S}_{c}$ and apply the corresponding transformation $\mathcal{T}_{\theta}^{s}$ during generation. This ensures that the output distribution remains uniformly balanced with respect to $\mathcal{C}$. For safe generation, we aggregate the transformations across all concepts in $\mathcal{S}_{\text{safe}}$, which includes violence and nudity since they broadly cover other safety concerns, such as shocking, harassment, and other harmful visuals. These transformations specifically correspond to “anti-violence” and “anti-sexual”, learned through negative prompting with respective target prompts. The aggregation is performed by computing the summation $\sum_{s\in\mathcal{S}_{\text{safe}}}\mathcal{T}_{\theta}^{s}(\bm{h})$. This aggregated transformation is then used to generate images that mitigate inappropriate content while preserving semantic fidelity during inference.
5 Experiments
We evaluate the effectiveness of the learned responsible concepts in promoting fair and safe image generation. All experiments are conducted using Stable Diffusion v1.4 (Rombach et al., 2022) and Stable Diffusion XL (Podell et al., 2024) to assess the efficacy of our approach. Note that the boldfaced values indicate the best results, while the underlined values represent the second-best results across all evaluations.
5.1 Fair Generation
Evaluation setting: We evaluate our method on the Winobias benchmark (Zhao et al., 2018), following the approaches in (Gandikota et al., 2024; Li et al., 2024; Orgad et al., 2023), which includes 36 professions with known gender biases. We learn transformations as constant functions linearly added to bottleneck activations, optimizing over 5000 iterations with batch size 1, as in Section˜4.2. Unlike Gandikota et al. (2024) and Shen et al. (2024), our approach does not rely on profession-specific data during training. Instead, we utilize the prompt “a person” to learn generalized directions that are applicable across professions. For fair comparison, we adopt the experimental setup from Li et al. (2024) to evaluate fairness. Five prompts per profession are used, including templates like “A photo of a $\langle\text{profession}\rangle$”. We also extend our evaluation to the Gender+ and Race+ Winobias datasets (Li et al., 2024), which introduce terms like “successful” to trigger stereotypical biases (Gandikota et al., 2024). Additional dataset details are in Section˜7.4.1 of Appendix.
Metrics: We perform quantitative and qualitative analysis to evaluate the performance of our proposed approach. We employ the modified average deviation ratio (DevRat), as defined in Li et al. (2024) to quantify the fairness of the generated images. We also measure the alignment of the generated images with the Winobias prompts (WinoAlign). Additionally, we assess image fidelity using the FID score (Heusel et al., 2017) on the COCO-30K validation set (FID (30K)), while image-text alignment is measured with the CLIP score (Radford et al., 2021a) using COCO-30K prompts under fair transformations (CLIP (30K)). Further details are provided in Section˜7.4.6 of Appendix.
Figure 2: Comparison of RespoDiff and SD in generating profession images by gender (top: women in first 4 columns, men in the rest) and race (bottom: Black in first 2, Asian in next 3, White in last 2). RespoDiff better reflects target attributes while maintaining fidelity to SD outputs.
Table 1: Comparison of gender fairness, alignment and quality across with SD v1.4.
Approach DevRat ($\downarrow$) WinoAlign ($\uparrow$) FID ($\downarrow$) CLIP ($\uparrow$)
SD (CVPR, 2022) 0.68 27.51 14.09 31.33
SDisc (CVPR, 2024) 0.17 26.61 23.59 29.94
FDF (ICLR, 2024) 0.40 23.90 15.22 30.63
BAct (CVPR, 2024) 0.57 27.67 17.07 30.54
RespoDiff 0.14 27.30 14.91 30.67
Table 2: Comparison of race fairness, alignment and quality across with SD v1.4.
Approach DevRat ($\downarrow$) WinoAlign ($\uparrow$) FID ($\downarrow$) CLIP ($\uparrow$)
SD (CVPR, 2022) 0.56 27.51 14.09 31.33
SDisc (CVPR, 2024) 0.23 26.80 17.47 30.27
FDF (ICLR, 2024) 0.32 23.15 14.94 30.59
BAct (CVPR, 2024) 0.45 27.63 17.20 30.47
RespoDiff 0.16 27.53 12.82 31.02
Results: We compare our approach against baselines including Stable Diffusion (SD) (Rombach et al., 2022), FDF (Shen et al., 2024), SDisc (Li et al., 2024) and BAct (Parihar et al., 2024) with additional details in Section˜7.4.2. Tables˜2 and 2 present a comparison of our approach to various baseline methods, focusing on various metrics across both gender and race biases. We further analyze extended biases within these categories, with additional comparisons provided in Section˜7.4.7. RespoDiff consistently achieves the lowest average deviation ratio in both gender and race biases, even in challenging settings, highlighting its superior performance in mitigating biases across different professions. As observed in Section˜7.4.8 of the Appendix, our method effectively eliminates gender and racial biases in a range of professions compared to Stable Diffusion. Although FDF performs better in certain professions like Secretary, likely due to training on profession-specific images, our approach improves fairness across all professions on average without being explicitly trained on profession-specific prompts. This highlights our model’s strong generalization ability across different professions. Additionally, RespoDiff remains robust to bias-inducing prompts with phrases such as ‘successful’, as detailed in Section˜7.4.7. We also provide additional comparisons in Section˜7.4.3.
Our approach achieves strong text-to-image alignment (WinoAlign) while maintaining the most balanced gender representation across fair concepts. Effective debiasing should preserve image fidelity and text alignment, particularly for non-stereotypical prompts which is evaluated using compute FID (30K) and CLIP (30K). As shown in Tables˜2 and 2, RespoDiff maintains image quality comparable to Stable Diffusion across gender and race debiased models while ensuring strong text-to-image alignment. Unlike SDisc, which operates in the bottleneck space but struggles with image quality, our approach effectively balances fairness without compromising generation quality.
Qualitative analyses in Fig.˜2 further support our quantitative findings. RespoDiff effectively modifies Stable Diffusion outputs to align with sensitive subconcepts like “woman”, “black race” etc. using learned transformations while preserving profession-specific attributes. Additional results are available in Section˜7.8. We also present additional qualitative comparisons of our approach with hard prompting and SDisc in Section˜7.4.4 and Section˜7.4.5, respectively.
5.2 Safe Generation
Evaluation setting: The training process is conducted for 1500 iterations with a batch size of 1 in the safe generation experiments. During evaluation, we generate images using prompts from the I2P benchmark (Schramowski et al., 2023), which consists of 4703 inappropriate prompts categorized into seven classes, including hate, shocking content, violence, and others.
Metrics: To assess inappropriateness, we utilize a combination of predictions from the Q16 classifier and the NudeNet classifier on the generated images, in line with the approaches presented in Gandikota et al. (2023); Schramowski et al. (2023); Li et al. (2024). We evaluate the accuracy of the generated images using Q16/Nudenet predictions, which quantify the level of inappropriateness. We also compute the FID and CLIP scores to assess image fidelity and image-text alignment using the COCO-30K prompts. We compare our approach against baselines such as SD, SDisc, ESD (Gandikota et al., 2023) and SLD (Schramowski et al., 2023). Further details are provided in Section˜7.5.2.
Figure 3: Qualitative comparison of safe generation. RespoDiff removes nudity and violenCE present in SD outputs, producing safer and more appropriate images.
Results: Table˜3 compares average Q16/NudeNet accuracies across all seven I2P benchmark classes, where our approach surpasses baselines by a 20% margin.
Table 3: Comparison of safety and image quality metrics across approaches with SD v1.4.
Approach I2P ($\downarrow$) FID (30K) ($\downarrow$) CLIP (30K) ($\uparrow$)
SD Rombach et al. (2022) 0.27 14.09 31.33
SDisc Li et al. (2024) 0.27 15.98 31.03
SLD Schramowski et al. (2023) 0.20 18.76 29.75
ESD Gandikota et al. (2023) 0.32 13.68 30.43
RespoDiff 0.16 17.89 31.10
Despite training on a limited safety set $\mathcal{S}_{\text{safe}}=\{\text{violence},\text{nudity}\}$, RespoDiff generalizes well to other safety-critical categories, as detailed in Section˜7.5.3. Qualitative results are shown in Fig.˜3. We further assess FID and CLIP scores, demonstrating that our approach maintains image quality comparable to Stable Diffusion on COCO-30K while achieving superior image-text alignment. Notably, our method not only enhances safety beyond SLD, which is designed specifically for safety, but also outperforms it in image-text alignment. While ESD and SDisc achieve higher image quality, our approach effectively filters inappropriate content without significant degradation in visual fidelity. Additional results are provided in Section˜7.8.
5.3 Responsible Generation using SDXL
Table 4: Comparison of gender fairness, alignment and quality metrics with SDXL.
Approach DevRat ($\downarrow$) WinoAlign ($\uparrow$) FID ($\downarrow$) CLIP ($\uparrow$)
SD 0.72 28.50 13.68 32.19
RespoDiff 0.26 28.02 14.63 32.03
Table 5: Comparison of race fairness, alignment and quality metrics with SDXL.
Approach DevRat ($\downarrow$) WinoAlign ($\uparrow$) FID ($\downarrow$) CLIP ($\uparrow$)
SD 0.57 28.53 13.68 32.19
RespoDiff 0.23 28.59 14.72 32.04
Figure 4: Comparison of RespoDiff and SDXL for the Doctor profession, steered towards: Woman, Black, Asian and White (left to right).
In this section, we explore the applicability of RespoDiff to large-scale generative models, specifically SDXL, for responsible T2I generation. We train our method using SDXL for the fair generation task and evaluate its effectiveness by comparing the results against the baseline SDXL model. The quantitative results are reported in Table˜4 and Table˜5.
As shown in Table˜4 and Table˜5, RespoDiff significantly mitigates biases related to gender and race that are present in SDXL, achieving substantial improvements. Importantly, our method preserves both the visual quality and image-text alignment of the original SDXL model. We also provide qualitative results demonstrating the transformation of concepts such as woman, Black race, Asian race, and White race in Fig.˜4 and Section˜7.8 of the Appendix. Notably, RespoDiff accurately steer generated images toward target concepts while maintaining consistency with the original content.
5.4 Ablations
We perform ablation studies to analyze RespoDiff, examining the impact of individual modules and the benefit of employing separate modules for responsible concept alignment and semantic alignment. We also assess the influence of different architectures for these modules and impact of using shared transformation on the overall performance in Section˜7.6 and Section˜7.7 of the Appendix.
Table 6: Effect of different modules used in our approach.
Approach DevRat ($\downarrow$) WinoAlign ($\uparrow$) FID (30K) ($\downarrow$) CLIP (30K) ($\uparrow$)
RAM 0.12 26.12 15.63 29.93
RAM + SAM (RespoDiff) 0.14 27.30 14.91 30.67
Effect of individual modules: We analyze the impact of the RAM and SAM modules on fairness and image quality. The results are summarized in Table˜6. RAM alone yields a lower deviation ratio, indicating reduced bias, but slightly compromises alignment (lower CLIP and WinoAlign scores). Adding SAM increases the deviation ratio but improves both image quality and alignment, highlighting its role in balancing fairness with generation fidelity.
6 Conclusion
Our proposed framework, RespoDiff, presents a significant advancement in responsible text-to-image generation by effectively ensuring fairness and safety of content generation. By leveraging a dual-module bottleneck transformation and a novel score-matching objective, our approach ensures responsible generation without degrading image quality. Extensive evaluations demonstrate its superior performance over existing methods, achieving a substantial improvement in generating fair and safe images across diverse prompts. Furthermore, integration of RespoDiff into large-scale models like SDXL highlights its practical applicability and scalability. This work represents a crucial step toward the development of more reliable T2I generative models, paving the way for safer AI-driven creativity.